Mathematics 242 – Principles of Analysis Definitions Quiz 6 – March 21, 2014

Your Name:

B) (5) Define: lim_{x→a} f(x) = L.

A) (5) Give an example of a function which is defined at x = 0 but where lim_{x→0} f(x) does not exist.